OpenStudy (anonymous):

Please help will fan and medal :( What is the solution set of 4x = y and 2x2 − y = 0? {(0.5, -0.5), (2, -8)} {(0, 0), (2, 8)} {(1.5, 2), (8, 2)} {(-1.5, -2), (2, 8)}

OpenStudy (anonymous):

@dan815 @Owlcoffee

OpenStudy (sleepyjess):

Hello! Do you know how to use the substitution method?

OpenStudy (anonymous):

Like plugging them in?

OpenStudy (sleepyjess):

Yes! We know what y is (4x), so let's plug that in to the second equation :)

OpenStudy (anonymous):

Erm, how so ._.,

OpenStudy (sleepyjess):

Is that supposed to be 2x^2?

OpenStudy (sleepyjess):

Okay, so you see that y = 4x right? the first equation

OpenStudy (anonymous):

Yes 2x^2 and I see it

OpenStudy (sleepyjess):

Okay, so instead of y in the second equation, we're going to put in 4x in place of that y

OpenStudy (anonymous):

Ahh okay. so then from there where would we go?

OpenStudy (sleepyjess):

Okay, so now we have, \(2x^2 -4y=0\), right? Let me make sure I'm right on the next step, one minute please :)

OpenStudy (sleepyjess):

*4x

OpenStudy (sleepyjess):

Okay, so do you know what the quadratic formula is?

OpenStudy (anonymous):

ax+bx=c^2?

OpenStudy (sleepyjess):

Actually, we have a^2 + bx + c = 0, c would be 0, we need to put those values into the quadratic formula (this may look scary but it's really not), \(x = \dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (sleepyjess):

and because c is 0, that makes it even easier :)

OpenStudy (anonymous):

Okay one second while I try ._.

OpenStudy (anonymous):

Uhm, I'm a bit stuck ._. @sleepyjess

OpenStudy (sleepyjess):

Okay, well we know a is 2, and b is -4, c is 0 because there is no c in our equation :) x = \(\dfrac{4\pm\sqrt{-4^2-4(2)(0)}}{2(2)}\) and now, we just work it all out

OpenStudy (sleepyjess):

Okay, now I'm questioning myself...

OpenStudy (jadzia):

4x = y and 2x2 − y = 0 let me just do it step-by-step 1. sub y= 4x into the 2nd equation: \(\sf 2x^2-4x=0\) 2. Solve for 'x' by quadratic formula (a=4, b=-4, c=0) or by factoring (which is much easier) I think jess can continue it from here now ^_^

OpenStudy (anonymous):

(Sprry my computer froze)

OpenStudy (jadzia):

\(\sf x= \dfrac{4\pm\sqrt{(-4)^2-4(2)(0)}}{2(2)}\)

OpenStudy (jadzia):

\(\sf x=\dfrac{4\pm\sqrt{16}}{4}\)

OpenStudy (jadzia):

now what is square root of 16? @WhatEven

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

but none of the options are 4 @Jadzia

OpenStudy (jadzia):

right..we are not done yet. we are only working on the radical part so now it will be like this: \(\sf x=\dfrac{4\pm 4}{4}\) which we will split into two parts: \(\sf x=\dfrac{4+ 4}{4}\) and \(\sf x=\dfrac{4- 4}{4}\) now what values are you going to get for x if you solve those two?

OpenStudy (anonymous):

2 and -2?

OpenStudy (jadzia):

2 is right, -2 is not... what is \(\sf x=\dfrac{4-4}{4}\)?

OpenStudy (jadzia):

@whateven still there?

OpenStudy (jadzia):

after you find the value of x on the second one.. the next step is to plug those x values on to the equation to solve for their corresponding y-components.

OpenStudy (anonymous):

Computer froze again ;/

OpenStudy (jadzia):

try restarting your computer

OpenStudy (anonymous):

I will if it does it again but it glitches and takes an hour to do so sometimes so I'd rather not risk that yet. I got 0

OpenStudy (anonymous):

@Jadzia still there?

OpenStudy (jadzia):

yes that's right. So from your choices, which them gives you 2 and 0 as x-coordinates?

OpenStudy (jadzia):

**of